OpenStudy (ebonyserrano):

A car travels 234 miles in 4 hours and 30 minutes. How many miles does it travel per hour?

OpenStudy (raffle_snaffle):

r*t = d, Where r is the rate, t is the time and d is the distance.

OpenStudy (raffle_snaffle):

t = 4.5 hours d = 234 miles Plug and play.

OpenStudy (colvyn01):

58.5

OpenStudy (colvyn01):

58.5 per hour

OpenStudy (raffle_snaffle):

close, but not the answer.

jhonyy9 (jhonyy9):

4 hours = 4*60 = 240 minutes 4 h 30 minutes = 240"30 = 270 minutes 234 miles ------ 270 min. x miles ------- 60 min. -------------------------- x = 60*234/270 = 52 miles so from this result that this car travels 52 miles per hour hope helped